A 1 kg ball and a 10 kg ball are dropped from a height of 10 m at the same time. In the absence of air resistance, O the 10 kg ball will take 10 times the amount of time to reach the ground. O the two balls will hit the ground at the same time. 0 there is not enough information to determine which ball will hit the ground first. O the 10 kg ball will hit the ground first. O the 1 kg ball will hit the ground first.
Answer #1

the two balls will hit the ground at the same time is the correct option.

By using the third equation of motion,

S = ut + (1/2)gt^2

The displacement is same for the two stones and initial velocity is zero for both and acceleration due to gravity is also same.

Let displacement for first stone and second stone be S1 and S2 respectively.

And time taken for first and second stone be t1 and t2 respectively

According to question,

S1 = S2

0 + (1/2)gt1^2 = 0 + (1/2)gt2^2

t1 = t2